3. Evaluate the following payoffs for the game pictured here: (a) ui(οι, Γ) for σ,-(ja, 1/4, 1/4, 1/4) (b) 112(oi, О) for σ,-(1/8, î/4, 1/43/8) 2 OA | 2,2 | 2,2 2,2 2, IA 4,2 1,3 IB 3,4 1,3
Answer #1

a)

P1 is playing mixed strategy σ1 = (1/4, 1/4, 1/4, 1/4)

P2 is playing column I

small u_1 (sigma_1, I) = rac{1}{4}*2 + rac{1}{4}*2 + rac{1}{4}*4 + rac{1}{4}* 3 = 11/4

b)

P1 is playing mixed strategy σι = (1/8. 1/4, 1/43/8)

P2 is playing column O

3 21

c)

P1 is playing mixed strategy σ1 = (1/4, 1/4, 1/4, 1/4)

P2 is playing column ơ-= (1 /3, 23)

21 1

2 2342223 1212 12 1212 12 12 12 12

d)

P1 is playing mixed strategy small sigma_1 = (0, 1/3, 1/6, 1/2)

P2 is playing column small sigma_2 = (2/3, 1/3)

3-33-63-23

2 1124 4 9 18 6 18 3